OpenStudy (anonymous):

4x+6y=12*62 x+y=12 18=x+y 9=y−2x Using the two equations above, write and solve the systems to find x and y.

OpenStudy (anonymous):

y - 2x = 9 y + x = 18 subtract -3x = 27 x = -9 sup this into first equn.: -9 + y = 18 y = 27 (x,y) = (-9,27)